# Energy landscapes of small SK spin glasses

### v1

@@ 4.1 Invariance classes and energy spectra for $N$=5
4.2 Invariance classes and energy spectra for $N$=6
@@
For $N=6$ the number of frustrated triangles is already not enough to distinguish the different invariance classes, because, for example, in classes 4, 5 and 6 this number is the same 8, so we have to determine the number of longer cycles to see the difference between them. The number of quadrangles is still not enough, because it coincides between classes 5 and 6. We have to go to counting also the pentagons in order to resolve these coincidences.
@@
[…]
@@ 5 Energy maps
In this Section we present a sample of energy maps for different $N$’s. In Subsection 5.1 we focus on the ground states and adjacent excited states to illustrate how the degenerate ground states begin to organize, already for small $N$’s, into clusters anticipating some elementary features of the ultrametric structure. In view of the large number of cases we will not go into details in each class, only in those that present some interesting properties. In Subsection 5.2 we make a few comments on higher energy minima and other isolated clusters and in Subsection 5.3 on the distribution of the ground state energies.
@@ 5.1.1 The case of $N$ = 3
Let us begin with the simplest possible case with $N=3$. There are just two classes here: the one with all couplings positive (a pure ferromagnet) and its gauge transformed variants, and the one with all couplings negative (the pure antiferromagnet) again with its gauge transforms. It is sufficient to consider a single representative of each class. For $N=3$ the phase space is just the ordinary cube which can be easily drawn and all 8 spin states can be shown in Fig. 4. The vertices in this graph correspond to the spin states; the edges connect neighboring spin states that differ only in the sign of a single spin.
[…]
@@
The left figure in Fig. 4 shows the phase space map of the $N=3$ ferromagnet. There are two degenerate ground states of energy $\epsilon_{0}\leavevmode\nobreak\ =-3$ and a six-fold degenerate excited state with energy $\epsilon_{1}\leavevmode\nobreak\ =+1$. The right figure in Fig. 4 is the map of the states of the antiferromagnet. This map is the reverse of that of the ferromagnet: there is a six-fold degenerate ground state ($\epsilon_{0}\leavevmode\nobreak\ =-1$) and a two-fold degenerate excited state ($\epsilon_{1}\leavevmode\nobreak\ =+3$). The antiferromagnetic ground states are all connected with each other into a network. This is a special case of the general feature of antiferromagnetic ground states for odd $N$: as can be easily seen, the ground state energy is $\epsilon_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ -\frac{1}{2}(N-1)$ and the degeneracy of the ground state is $\mu_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ \binom{N}{\frac{N+1}{2}}+\binom{N}{\frac{N-1}{2}}$. These ground states are connected with each other with $\frac{N+1}{2}$ links. The remaining $\frac{N-1}{2}$ couplings connect the ground states to the first excited states.
The ferromagnetic case is just the opposite of the antiferromagnetic one, it is the highest excited states here that are the most numerous and connected with each other with $\frac{N+1}{2}$ links, with an occasional path leading down toward the ground state. This arrangement can be called a plateau with some escape routes. If we think of an optimization task with such a phase space and initiate the search on the plateau, our program may spend a long time wandering around on the top before finding a way to descend.
@@ 5.1.2 The case of $N$ = 4
The left Fig. 5a shows the $N=4$ ferromagnet, Class 1 in Table 1. This map corresponds to our usual way of speaking about an Ising ferromagnet: the paths starting from one of the ground states lead through the highest energy states before arriving at the ”other half of phase space” and eventually at the other ground state. (The meaning of the two halves of phase space is the following: spin vectors whose components sum to a positive value belong to one half of the phase space, those with a negative sum of their components to the other. For even $N$ there are vectors whose components sum to zero; these lie in the transition zone between the two halves.)
@@
This state of affairs changes when we consider the fully frustrated configuration whose map is given in Fig. 5c. The ground states are now 6-fold degenerate (Class 3 in Table 1) and can be reached from each other by going through the first excited states, so the communication between the regions of phase space is much easier and the distinction between its two halves is nearly meaningless. (Of course, this is true only as long as there is no external field present, but that is beyond the scope of the present paper.)
This is leading to an easy generalization again. For a general even $N$ the ground state energy of the pure antiferromagnetic configuration (and its gauge and permutation equivalents) is $\epsilon_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ -N/2$ and its degeneracy is $\mu_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ \binom{N}{\frac{N}{2}}$. Every ground state has $N$ next neighbors among the first excited states and do not have other neighbors apart from those.
@@ 5.1.3 Energy maps for $N$=5
[…]
Figure 6: Phase space map of three classes for $N=5$. Only the ground states (black) and some of the 1st excited states (blue) are shown, with (a) corresponding to Class 3, (b) to Class 6, and (c) to Class 4 in Table 2.
@@
It may be surprising to see such a highly symmetric map arising from the asymmetric distribution of positive and negative links shown in Fig. 2.4. We have to realize, however, that the representation in that figure is just the most economic one, in that it contains the least number of negative links. Through a few steps of gauge transformations and permutations one can show that an equivalent representation of this class can be given either by a pentagon or a five-pointed star of negative links, the others being positive, as shown in Fig. 7. This explains the five-fold symmetry of the energy map. It is also a demonstration of how highly symmetric energy landscapes can be hidden behind the curtain of gauge transformations.
[…]
@@
These members of Class 4 display some further interesting properties. If we change the signs of each bond in Fig. 7a, we get Fig. 7b, and vice versa. So the energy spectrum of either one of them is the same as that of the negative of the other, while on the other hand they are gauge and permutation equivalent, so have the same spectrum. Therefore this spectrum must be symmetric w.r.t. the origin, which indeed it is, see Class 4 in Table 2. The two representatives of Class 4 shown in Fig. 7 are then examples of sign-symmetric signed graphs. It may be interesting to identify classes with a similar property for higher $N$ values.
[…]
@@
5.1.4 Energy maps for $N$=6
The first few classes do not show remarkable features: we have doubly degenerate ground states which are separated from each other by a barrier, which is becoming lower and lower as we add more negative bonds. This picture changes in Class 6, 9, and 11 where we can observe how the ground states start to divide the phase space among themselves. The corresponding energy maps are shown in Fig. 8. As can be seen, the ground states have their small basins of attraction separated by low, one- or two energy unit barriers. At this small $N$ one could not expect to find deep energy valleys, but it is not hard to imagine that with increasing $N$ the fragmentation of phase space continues and the barriers grow. In Class 9 and 11 we can also see that the ground states are organized in small families, within which the barriers are lower than between them. We regard these cases as the appearance of the fragmentation of phase space in an embryonic form. The fast growing number of cases make it very hard to analyze the classes to a comparable detail for larger $N$’s, but we are sure that with $N$ growing more and more branches of the ultrametric tree would show up with higher and higher barriers between them.
@@ 5.2 Local minima and isolated clusters for energies higher the the ground state
Our findings are the following. There are no local minima or isolated clusters up to and including $N=5$. For $N=6$ we find three classes with 2, 4 resp. 8 isolated local minima: single spin states whose surrounding states are one step higher in energy. For $N=7$ there is no local minimum or isolated cluster above the ground state(s).
Things become interesting at $N=8$ where different types of isolated cluster start to develop. Beside the single-point local minima, more extended ones appear. We have identified 55 classes with extended clusters. Of these 6 classes have a single large extended cluster consisting of 60, 42, 28, 24, 24 and 16 spin states, respectively. In 45 classes we found two mirror symmetric clusters the largest ones containing $2\times 18$, $2\times 17$, in three classes $2\times 12$ and in 4 classes $2\times 10$ states. In addition, we found 4 classes having 4 clusters each, with sizes $4\times 4$ in one of the classes, and $4\times 2$ in three cases.
@@
The full display of all these isolated clusters would take up too much space, so we present here just the most interesting one. This cluster consist of 60 alternating first and second excited states which are sitting on the surface of a torus! The rectangle in Fig. 10 is a planar representation of this torus where periodic boundary conditions are meant. The spin states neighboring the ones on the surface of the torus have one or two unit higher energy. The torus goes around in phase space, but the states belonging to it do not exhaust the whole set of first and second excited states, so one can walk around it when going from the ground states to the higher energy states without necessarily stepping into this shallow minimum.
For $N=9$ there are 2 classes with clusters of size 2 (in one of them there are 4 such clusters, in the other 2 clusters), and there are single-state local minima in 75 classes. These numbers are to be compared with the total number of classes 1625 at $N=9$.
Note the marked difference between the even and odd $N$ cases: the even $N$ classes tend to be richer in detail. For higher $N$’s this difference presumably disappears.
@@
A final remark on the extended objects appearing in these energy maps: The icosahedron popping up in one of the ground state maps at $N=6$ or the torus at $N=8$ are constituted by a set of spin vectors, that is some vertices of a 6-, resp. 8-dimensional cube. The role of the Hamiltonian is merely to designate the spin vectors that belong to a given energy. The icosahedron, the torus and all the other objects exhibited in this paper are simply the intersections of the given energy surface with the hypercube, but these objects are there in the cube, waiting to be picked out.
@@
5.3 Towards self-averaging
If we throw all the $2^{N(N-1)/2}$ different SK samples into an urn and draw a few at random, we will find that for large $N$ the ground state energy of the selected samples will be almost always the same: the probability of drawing this particular value of the ground state energy will be concentrated on a subset of samples which dominate the statistics. This phenomenon is called ”self-averaging” and is the justification of averaging over the random interactions in the limit of large $N$. Since all the the gauge classes have the the same number of members, it is enough, in fact, to throw one representative sample of each gauge class into the urn and test the distribution of the ground states over the permutation classes only.
The difference between the energy of the antiferromagnetic ground state and the ferromagnetic one is $\epsilon^{anti}_{0}-\epsilon^{ferro}_{0}=\frac{1}{4}N(N-2)$ for $N$ even, and $\frac{1}{4}(N-1)^{2}$ for $N$ odd. The energy steps between adjacent configurations is 2 units, so the number of steps is $K^{even}_{steps}=\frac{1}{4}N(N-2)+1$ for $N$ even, and $K^{odd}_{steps}=\frac{1}{4}(N-1)^{2}+1$ for $N$ odd. The total number of permutation configurations is $2^{(N-1)(N-2)/2}$ which agrees with the number of steps for $N=3$, but larger than that for any $N>3$. Therefore, for $N>3$ more than one energy must occupy some of the steps, by the pigeon hole principle. For large $N$’s the number of pigeons will vastly exceed the number of holes, so on some steps we will find a very large number of configurations with the same ground state energy. The distribution of ground state energies is shown in Fig. 11. We can see, how the distribution is becoming sharper and sharper with $N$ growing.
Figure 11: Multiplicity of ground states, normalized to fit the figure, for $N=6,\ldots,9$.
6 Conclusion
We have investigated all the SK spin glass samples up to $N=9$ vertices and classified them into invariance classes according to the gauge and permutation symmetry. In the course of this work we made some contact with the theory of signed graphs. We have the impression that despite the evident overlap between the problem of signed graphs and that of spin glasses not all the mutual advantages have been exploited. For example, it would be interesting to know which are the properties of the interaction matrix $J$ that determine the fragmentation of phase space, or the emergence of such surprising structures as e.g. the torus in one of the $N=8$ classes. There is virtually no discussion of spin glasses without invoking the picture of a rugged energy landscape, but we have never heard such a feature as the torus being mentioned.
It is obvious that the invariance class determines the spin glass spectrum, but we do not know whether the reverse statement is true. All we could ascertain was that up to $N=9$ the spin glass energy spectra reliably identified the different classes even in the exceptional case of the two $N=8$ matrices where the number of frustrated cycles failed to do so.
It is often pointed out that spin glasses possess equilibrium states, or ground states that are in no obvious symmetry relation with each other. This is certainly true if one draws the interactions from a continuous distribution and averages over them. It is nothing short of a miracle that the ultrametric symmetry shines through even after all this. We used discrete couplings and small sizes with no averaging. In this setup it is natural that the energy maps we found reflect the symmetries of the interaction matrix. What can sometimes obfuscate the picture is how these symmetries may be hidden among the gauge invariant versions of the interaction.
@@ 6 Conclusion
In a way, there is nothing extraordinary about this: we are dealing with a non-convex energy landscape. What is remarkable is that such a toy model as the SK Hamiltonian is able to reflect a universal feature of Nature. Whether we look at the atomic nuclei, the atoms, molecules, living beings, or ecosystems we see the same pattern: a few sorts of building blocks with competing interactions arrange themselves into a number of stable structures that act as attractors for the unstable and metastable ones. But the essential point is that these attractors are, as a rule, not unique. The picture we have found in our small spin glass samples is the appearance of multiple equilibria [19] in an embryonic form. The main message is how soon these equilibria appear and how soon they start to arrange themselves into families.

### v3

@@ 4.1 Invariance classes and energy spectra for NN=5
4.2 Invariance classes and energy spectra for NN=6
@@
For $N=6$ the number of frustrated triangles is already not enough to distinguish the different invariance classes, because, for example, in classes 4, 5 and 6 this number is the same 8, so we have to determine the number of longer cycles to see the difference between them. The number of quadrangles is still not enough, because it coincides between classes 5 and 6. We also have to count the pentagons in order to resolve these coincidences.
@@
We have determined the isospectral classes for $N=7,\penalty 10000\ 8$ and $9$ as well. Their numbers are 54, 219, and 1625, respectively. We also have the number of frustrated cycles and the energy spectra for these cases. Such a large amount of data cannot be squeezed into this paper, therefore we placed it on the web site [12].
@@ 5 Energy maps
In this Section we present a sample of energy maps for different $N$’s. In Subsection 5.1 we focus on the ground states and adjacent excited states to illustrate how the degenerate ground states begin to organize, already for small $N$’s, into clusters anticipating some elementary features of the ultrametric structure. In view of the large number of cases we will not go into details in each class, only in those that present some interesting properties. In Subsection 5.2 we make a few comments on higher energy minima and other isolated clusters.
@@ 5.1.1 The case of $N$ = 3
Let us begin with the simplest possible case with $N=3$. There are just two classes here: the one with all couplings positive (a pure ferromagnet) and its gauge transformed variants, and the one with all couplings negative (the pure antiferromagnet) again with its gauge transforms. It is sufficient to consider a single representative of each class. For $N=3$ the phase space is just the ordinary cube that is easy to draw with all 8 spin states shown as in Fig. 4. The vertices in this graph correspond to the spin states; the edges connect neighboring spin states that differ only in the sign of a single spin.
a
b
@@
The figure on the left in Fig. 4 shows the phase space map of the $N=3$ ferromagnet. There are two degenerate ground states of energy $\epsilon_{0}\penalty 10000\ =-3$ and a six-fold degenerate excited state with energy $\epsilon_{1}\penalty 10000\ =+1$. The figure on the right in Fig. 4 is the map of the states of the antiferromagnet. This map is the reverse of that of the ferromagnet: there is a six-fold degenerate ground state ($\epsilon_{0}\penalty 10000\ =-1$) and a two-fold degenerate excited state ($\epsilon_{1}\penalty 10000\ =+3$). The antiferromagnetic ground states are all connected with each other in a network. This is a special case of the general property of antiferromagnetic ground states for odd $N$: as can easily be seen, the ground state energy is $\epsilon_{0}\penalty 10000\ =\penalty 10000\ -\frac{1}{2}(N-1)$ and the degeneracy of the ground state is $\mu_{0}\penalty 10000\ =\penalty 10000\ \binom{N}{\frac{N+1}{2}}+\binom{N}{\frac{N-1}{2}}$. These ground states are connected to each other by $\frac{N+1}{2}$ links. The remaining $\frac{N-1}{2}$ couplings connect the ground states to the first excited states.
The ferromagnetic case is just the opposite of the antiferromagnetic one, it is the highest excited states here that are the most numerous and connected with each other by $\frac{N+1}{2}$ links, with an occasional path leading down toward the ground state. This arrangement can be called a plateau with some escape routes. If we think of an optimization task with such a phase space and initiate the search on the plateau, our program may spend a long time wandering around at the top before finding a way to descend.
@@ 5.1.2 The case of $N$ = 4
The left Fig. 5a shows the $N=4$ ferromagnet, Class 1 in Table 1. This map corresponds to our usual way of speaking about an Ising ferromagnet: the paths starting from one of the ground states lead through the highest energy states before arriving at the ”other half of phase space” and eventually at the other ground state. (The meaning of the two halves of phase space is obvious: spin vectors whose components sum to a positive value belong to one half of the phase space, those with a negative sum of their components to the other. For even $N$ there are vectors whose components sum to zero; these lie in the transition zone between the two halves.)
@@
This state of affairs changes when we consider the fully frustrated configuration whose map is given in Fig. 5c. The ground states are now six-fold degenerate (Class 3 in Table 1) and can be reached from each other by going through the first excited states, so the communication between the regions of phase space is much easier and the distinction between its two halves is nearly meaningless. (Of course, this is true only as long as there is no external field present, but that is beyond the scope of the present paper.)
This leads us to an easy generalization again. For a general even $N$ the ground state energy of the pure antiferromagnetic configuration (and its gauge and permutation equivalents) is $\epsilon_{0}\penalty 10000\ =\penalty 10000\ -N/2$ and its degeneracy is $\mu_{0}\penalty 10000\ =\penalty 10000\ \binom{N}{\frac{N}{2}}$. Every ground state has $N$ next neighbors among the first excited states and does not have other neighbors apart from those.
@@ 5.1.3 Energy maps for NN=5
b
c
Figure 6: Phase space map of three classes for $N=5$. Only the ground states (black) and some of the first excited states (blue) are shown, with (a) corresponding to Class 3, (b) to Class 6, and (c) to Class 4 in Table 2.
@@
It may be surprising to see such a highly symmetric map arising from the asymmetric distribution of positive and negative links shown in Fig. 2. We have to realize, however, that the representation in that figure is just the most economic one, in that it contains the least number of negative links. Through a few steps of gauge transformations and permutations one can show that an equivalent representation of this class can be given either by a pentagon or a five-pointed star of negative links, the others being positive, as shown in Fig. 7. This explains the five-fold symmetry of the energy map. It is also a demonstration of how highly symmetric energy landscapes can be hidden behind the curtain of gauge transformations.
a
b
@@
These members of Class 4 display some further interesting properties. If we change the signs of each bond in Fig. 7a, we get Fig. 7b, and vice versa. So the energy spectrum of either one of them is the same as that of the negative of the other, while on the other hand they are gauge and permutation equivalent, so have the same spectrum. Therefore this spectrum must be symmetric to the origin, which it is indeed, see Class 4 in Table 2. The two representatives of Class 4 shown in Fig. 7 are then examples of sign-symmetric signed graphs. It may be interesting to identify configurations with a similar property for higher $N$ values.
Another remarkable property of this class is that the cube of these interaction matrices $J$ is proportional to the matrix itself: $J^{3}\penalty 10000\ =\penalty 10000\ 5J$.
@@
5.1.4 Energy maps for NN=6
The first few classes do not show remarkable features: we have doubly degenerate ground states which are separated from each other by a barrier, which is becoming lower and lower as we add more negative bonds. This picture changes in Class 6, 9, and 11 where we can observe how the ground states start to divide the phase space among themselves. The corresponding energy maps are shown in Fig. 8. As can be seen, the ground states have their small basins of attraction separated by low, one- or two energy unit barriers. At this small $N$ one could not expect to find deep energy valleys, but it is not hard to imagine that with increasing $N$ the fragmentation of phase space continues and the barriers grow. In Class 9 and 11 we can also see that the ground states are organized in small families, within which the barriers are lower than between them. We regard these cases as the appearance of the fragmentation of phase space in an embryonic form. The fast growing number of cases make it very hard to analyze the classes in a comparable detail for larger $N$’s, but we are sure that with $N$ growing more and more branches of the ultrametric tree would show up with higher and higher barriers between them.
@@ 5.2 Local minima and isolated clusters for energies higher the the ground state
Our findings are as follows. There are no local minima or isolated clusters up to and including $N=5$. For $N=6$ we find three classes with 2, 4 resp. 8 isolated local minima: single spin states whose surrounding states are one step higher in energy. For $N=7$ there is no local minimum or isolated cluster above the ground state(s).
Things become interesting at $N=8$ where different types of isolated cluster start to develop. In addition to the single-point local minima, more extended ones appear. We identified 55 classes with extended clusters. Of these 6 classes have a single large extended cluster consisting of 60, 42, 28, 24, 24 and 16 spin states, respectively. In 45 classes we found two mirror symmetric clusters the largest ones containing $2\times 18$, $2\times 17$, in three classes $2\times 12$ and in 4 classes $2\times 10$ states. In addition, we found 4 classes having 4 clusters each, with sizes $4\times 4$ in one of the classes, and $4\times 2$ in three cases.
@@
The full display of all these isolated clusters would take up too much space, so we present here just the most interesting one. This cluster consists of 60 alternating first and second excited states which are sitting on the surface of a torus! The rectangle in Fig. 10 is a planar representation of this torus where periodic boundary conditions are meant. The spin states neighboring the ones on the surface of the torus have one or two unit higher energy. The torus goes around in phase space, but the states belonging to it do not exhaust the whole set of first and second excited states, so one can walk around it when going from the ground states to the higher energy states without necessarily stepping into this shallow minimum.
For $N=9$ there are 2 classes with clusters of size 2 (in one of them there are 4 such clusters, in the other 2 clusters), and there are single-state local minima in 75 classes. These numbers are to be compared with the total number of classes 1625 for $N=9$.
Note the marked difference between the even and odd $N$ cases: the even $N$ classes tend to be richer in detail.
@@
A final remark on the extended objects appearing in these energy maps: The icosahedron popping up in one of the ground state maps at $N=6$ or the torus at $N=8$ are constituted by a set of spin vectors, that is some vertices of a 66-, resp. 88-dimensional cube. The role of the Hamiltonian is merely to designate the spin vectors that belong to a given energy. The icosahedron, the torus and all the other objects exhibited in this paper are simply the intersections of the given energy surface with the hypercube, but these objects are there in the cube, waiting to be picked out.
@@
6 Summary
We have investigated all the SK spin glass samples up to $N=9$ vertices and classified them into isospectral classes with the help of the gauge- and permutation symmetries. In the course of this work we made some contact with the theory of signed graphs. We have the impression that despite the evident overlap between the problem of signed graphs and that of spin glasses not all the mutual advantages have been fully exploited. For example, it would be interesting to know which are the properties of the interaction matrix $J$ that determine the fragmentation of phase space, or the emergence of such surprising structures like e.g. the torus in one of the $N=8$ classes. There is virtually no discussion of spin glasses without invoking the picture of a rugged energy landscape, but we have never heard such a feature as the torus being mentioned.
It is obvious that the invariance class determines the spin glass spectrum, but we do not know whether the reverse statement is true. We intend to return to this question in future work.
It is often pointed out that spin glasses possess equilibrium states, or ground states that are in no obvious symmetry relation with each other. This is certainly true if one draws the interactions from a continuous distribution and averages over them. It is nothing short of a miracle that the ultrametric symmetry shines through even after all this. We used discrete couplings and small sizes with no averaging. In this setup it is natural that the energy maps we find reflect the symmetries of the interaction matrix. What can sometimes obfuscate the picture is how these symmetries may be hidden among the gauge invariant versions of the interaction.
@@ 6 Summary
In a way, there is nothing extraordinary about this: we are dealing with a non-convex energy landscape. What is remarkable is that such a toy model as the SK Hamiltonian is able to reflect a universal feature of Nature. Whether we look at the atomic nuclei, the atoms, molecules, living beings, or ecosystems we see the same pattern: a few kinds of building blocks with competing interactions arrange themselves into a number of stable structures that act as attractors for the unstable and metastable ones. But the essential point is that these attractors are, as a rule, not unique. The picture we have found in our small spin glass samples is the appearance of multiple equilibria [20] in an embryonic form. The main message is how soon these equilibria appear and how soon they start to arrange themselves into families.